OpenStudy (nerdy_3000):

Solve the equation for y then substitute the value of x into the equation? -xy+3x=30 x=15

OpenStudy (anonymous):

First of all, you add xy both sides to get 3x=30+xy, then you subtract 30 from both sides to get 3x-30=xy,then y=3(x-10)/x, now just plug in the value of x which is 15 in this case.

OpenStudy (nerdy_3000):

Where did you get the 10?

OpenStudy (anonymous):

3x-30=3(x-10), by taking 3 common.

OpenStudy (anonymous):

If you have confusion about that, just take 3x-30. It would give the same answer.

OpenStudy (nerdy_3000):

Okay thank you! Can you help me with one more problem?

OpenStudy (nerdy_3000):

The area of a triangle a. solve the area formula for h What is the height of a triangle with an area of 20in^2 and base of 4in

OpenStudy (anonymous):

 Use the given relation, A=Area,b=Base,h=Height , you should get the value of h then. Try it for yourself.